# Improved belief propagation is sufficient for real-time decoding of quantum memory

### v1

@@ Abstract
Even on a CPU, OSD can be slow, prompting the development of faster follow-ups for BP. Ambiguity Clustering (AC) [28] and Localized Statistics Decoding (LSD) [29] are examples of relatively fast clustering-based decoders proposed for this purpose, which are similar to Union-Find generalized to qLDPC codes [30]. Unfortunately, while Union-Find has a compact FPGA implementation for surface codes, its qLDPC generalizations require expensive matrix operations, albeit for smaller matrices than BP+OSD. Without further research, it is therefore unclear if clustering-based decoders can be implemented compactly on FPGAs for practical qLDPC decoding. Recently, new search-based decoders [31, 32] and decimated-BP [33] have emerged which are not only faster than BP-OSD, but also achieve significantly lower logical error rates. However, these approaches also pose FPGA challenges due to their complex data structures.
@@
In this paper we introduce a new flexible qLDPC decoder called Relay-BP (“Relay-ensembling with locally-averaged memory”), which satisfies all the criteria in Table 1. Relay-BP is based on modifications to BP that retain its message-passing structure thereby ensuring a compact and fast FPGA implementation. We first build on prior work that introduced memory terms to dampen oscillations in BP [34, 35, 36], adopting a $\mathrm{GF}(2)$ version of EWAInit-BP [36], which we refer to as Mem-BP. Observing that varying memory strengths across node types improves convergence [37, 38], we generalize this into DMem-BP using highly disordered memory strength distributions. Secondly, a relay ensembling technique is incorporated in which successive DMem-BP runs are initialized with the previous run’s final marginals. Finally, to help escape from trapping sets, we allow the memory strengths to become negative. We illustrate these features in Figure 1, and provide more explicit definitions and analysis of Relay-BP in the rest of the text.
@@ I Decoding
A decoding problem can be formally defined  [^2] See the appendix for details or reference [31] by a check matrix $\mathbf{H}\in\mathbb{F}_{2}^{M\times N}$, an action matrix $\mathbf{A}\in\mathbb{F}_{2}^{K\times N}$ and a probability vector $\mathbf{p}=(p_{0},\dots,p_{N-1})$. We assume each error location $j\in\{0,\dots,N-1\}=[N]$ independently experiences an error with probability $p_{j}$. Let $\mathbf{e}\in\mathbb{F}_{2}^{N}$ be the bit-string representing the unknown compound error and let $\bm{\sigma}=\mathbf{H}\mathbf{e}\in\mathbb{F}_{2}^{M}$ be the observed syndrome. Our task is to infer a correction $\mathbf{\hat{e}}$ using BP such that $\mathbf{H}\bm{\hat{e}}=\bm{\sigma}$ and $\mathbf{A\hat{e}}=\mathbf{Ae}$. Different choices of $\mathbf{H}$, $\mathbf{A}$, and $\mathbf{p}$ define decoding problems across a wide range of settings: from classical codes to separate X and Z decoding for quantum CSS codes [40, 41], and further to fault-tolerant quantum circuits under correlated circuit noise. Let $G$ be the associated decoding graph where error nodes correspond to columns of $\mathbf{H}$ and check nodes correspond to rows of $\mathbf{H}$. We denote the neighbors of node $w$ in $G$ as $\mathcal{N}(w)$. A family of decoding problems is said to be qLDPC if the degree of $G$ does not grow with the problem size (we assume degree at most ten is desirable for practical message passing algorithms).
@@ III Relay-BP-$S$
Figure 2: Circuit-noise decoding examples. (Top) Relay-BP-1 logical error rate heatmaps at $p=3\times 10^{-3}$ vs. memory strength intervals. Circles mark intervals used for decoding: $[-0.24,0.66]$ (gross), $[-0.161,0.815]$ (two-gross), $[-0.25,0.85]$ (surface). The dashed line indicates a threshold above which negative $\gamma_{j}$ are present. (Bottom) Relay-BP outperforms BP+OSD+CS-10 on the gross and two-gross codes, and performs comparably to Matching on the surface code. Top panels share their width scale but have separate logical error rate heat maps; bottom panels share a logical error rate scale.
@@ IV Decoding examples
We consider two decoding strategies. XYZ-decoding directly computes a correction $\bm{\hat{e}}$ from $\mathbf{H}$ and $\bm{\sigma}$. XZ-decoding decomposes $\bm{\sigma}$ into $\bm{\sigma}_{X}$ and $\bm{\sigma}_{Z}$, which are independently decoded using the derived check matrices $\mathbf{H}_{X}$ and $\mathbf{H}_{Z}$ to obtain $\bm{\hat{e}}_{X}$ and $\bm{\hat{e}}_{Z}$. These partial corrections are then combined to infer $\bm{\hat{e}}$. XZ-decoding simplifies decoding by using smaller objects but may degrade performance by modeling X and Z errors as independent. All three decoding examples are CSS-type; we use XZ-decoding throughout unless explicitly stated otherwise. The XZ-decoding matrices have approximate dimensions 1k×$\times$9k, 2k×$\times$26k and 1k×$\times$7k for gross, two-gross and surface code examples respectively, with 2k×$\times$71k, 5k×$\times$213k and 1k×$\times$24k for the XYZ-decoding matrices.
In all Relay-BP simulations, we set the max iterations to $T_{r}=60$ for each leg, except the first, which uses $T_{0}=80$ due to slower initial convergence. The first leg uses a uniform memory strength of $\gamma=0.125$ for the gross and two-gross codes, and $\gamma=0.35$ for the surface code, based on preliminary sweep results. In all subsequent legs, each $\gamma_{j}$ is drawn independently and uniformly from a common range $[\gamma_{\text{center}}-\gamma_{\text{width}}/2,\gamma_{\text{center}}+\gamma_{\text{width}}/2]$.
@@ V Memory strength selection
Relay-BP’s accuracy can be improved significantly by appropriate problem-dependent choice of memory strengths. Figure 2 shows heatmaps for each of the three decoding problems at $p=3\times 10^{-3}$, reporting logical error rates of Relay-BP-1 ($R=301$) with memory strengths from intervals with various values of $\gamma_{\text{center}}$ and $\gamma_{\text{width}}$. Each tile uses up to 500,000 samples and most tiles are within 10% relative accuracy.
A striking feature in the heatmap for the gross and two-gross codes (and to a lesser extent for the surface code) is a performance hotspot above the diagonal, corresponding to intervals that include negative memory strengths. Intervals marked with circles in the heatmaps are used for all subsequent simulations.
@@ VII Real-time decoding
We assess the real-time feasibility of Relay-BP using XZ-decoding of the gross code at $p=3\times 10^{-3}$. This 12-cycle circuit at 1 $\mu$s per cycle can accommodate $\sim 600$ BP iterations to achieve real-time decoding without backlog. This estimate assumes a 20 ns iteration time, based on Ref. [27], which reports 16 ns for a slightly simpler decoding problem using a fully parallel FPGA implementation with dedicated resources for each variable and check node.
@@
In Figure 5 in the appendix, we observe that with 80 iterations, Relay-BP-1 achieves an order of magnitude lower logical error rate for XYZ-decoding than for XZ-decoding on the same problem. While we expect the FPGA iteration time only grows moderately with decoding matrix dimensions, we do expect XYZ-decoding ($\sim$2k×$\times$71k) to be somewhat slower than XZ-decoding ($\sim$1k×$\times$9k), reducing the effective iteration budget below the 600 estimated for XZ. For larger codes, the increased problem size will further raise iteration time and may exceed a single FPGA’s capacity if using the fully parallel implementation strategy of Ref. [27]. This work focuses on quantum memory, where avoiding the backlog problem requires a low average-case decoding time. Logical operations—especially those involving multiple code blocks—not only increase problem size but may also require decoding to complete before the next logical operation is selected, such that the tail of the decoding time distribution is also crucial [49].
@@ Appendix A Decoding objects for logical memory circuits
For the surface code data, we compute the logical error rate for X type failure, and multiply by two to account for Z type failure too.
@@ Appendix B Decoding
Let $\mathbf{p}\in\mathbb{F}_{2}^{N}$ be the probability vector where $p_{j}$ is the probability that error $j$ occurs. The discrete noise is then represented as an error vector $\mathbf{e}\in\mathbb{F}_{2}^{N}$, where we make the assumption that each bit is drawn independently according to a probability vector $\mathbf{p}$, such that:
@@ 9             $\hat{e}_{j}(t)\leftarrow\mathrm{HD}\bigl{(}M_{j}(t)\bigr{)}$;
10             if $\mathbf{H\hskip 0.83298pt\mathbf{\hat{e}}}(t)=\bm{\sigma}$ then

### v2

@@ Abstract
Even on a CPU, OSD can be slow, prompting the development of faster follow-ups for BP. Ambiguity Clustering (AC) [28] and Localized Statistics Decoding (LSD) [29] are examples of relatively fast clustering-based decoders proposed for this purpose, which are similar to Union-Find generalized to qLDPC codes [30]. Unfortunately, while Union-Find has a compact FPGA implementation for surface codes, its qLDPC generalizations require expensive matrix operations, albeit for smaller matrices than BP+OSD. Without further research, it is therefore unclear if clustering-based decoders can be implemented compactly on FPGAs for practical qLDPC decoding. Recently, new search-based decoders [31, 32] and decimated-BP [33] have emerged which are not only faster than BP-OSD, but also achieve significantly lower logical error rates. However, these approaches also pose FPGA challenges due to their complex data structures. Other attempts aimed at circumventing costly follow-ups for BP were unable to match these lower error rates [34] or so far only tested with simplified noise models  [35, 36].
@@
In this paper we introduce a new flexible qLDPC decoder called Relay-BP (“Relay-ensembling with locally-averaged memory”), which satisfies all the criteria in Table 1. Relay-BP is based on modifications to BP that retain its message-passing structure thereby ensuring a compact and fast FPGA implementation. We first build on prior work that introduced memory terms to dampen oscillations in BP [37, 38, 39, 40], adopting a $\mathrm{GF}(2)$ version of EWAInit-BP [40], which we refer to as Mem-BP. Observing that varying memory strengths across node types improves convergence, we generalize this into DMem-BP using highly disordered memory strength distributions. Secondly, a relay ensembling technique is incorporated in which successive DMem-BP runs are initialized with the previous run’s final marginals. Finally, to help escape from trapping sets, we allow the memory strengths to become negative. We illustrate these features in Figure 1, and provide more explicit definitions and analysis of Relay-BP in the rest of the text.
@@ I Decoding
A decoding problem can be formally defined  [^2] See the appendix for details or reference [31] by a check matrix $\mathbf{H}\in\mathbb{F}_{2}^{M\times N}$, an action matrix $\mathbf{A}\in\mathbb{F}_{2}^{K\times N}$ and a probability vector $\mathbf{p}=(p_{0},\dots,p_{N-1})$. We assume each error location $j\in\{0,\dots,N-1\}=[N]$ independently experiences an error with probability $p_{j}$. Let $\mathbf{e}\in\mathbb{F}_{2}^{N}$ be the bit-string representing the unknown compound error and let $\bm{\sigma}=\mathbf{H}\mathbf{e}\in\mathbb{F}_{2}^{M}$ be the observed syndrome. Our task is to infer a correction $\mathbf{\hat{e}}$ using BP such that $\mathbf{H}\mathbf{\hat{e}}=\bm{\sigma}$ and $\mathbf{A\hat{e}}=\mathbf{Ae}$. Different choices of $\mathbf{H}$, $\mathbf{A}$, and $\mathbf{p}$ define decoding problems across a wide range of settings: from classical codes to separate X and Z decoding for quantum CSS codes [42, 43], and further to fault-tolerant quantum circuits under correlated circuit noise. Let $G$ be the associated decoding graph where error nodes correspond to columns of $\mathbf{H}$ and check nodes correspond to rows of $\mathbf{H}$. We denote the neighbors of node $w$ in $G$ as $\mathcal{N}(w)$. A family of decoding problems is said to be qLDPC if the degree of $G$ does not grow with the problem size (we assume degree at most ten is desirable for practical message passing algorithms).
@@ III Relay-BP-$S$
Figure 2: Circuit-noise decoding examples. (Top) Relay-BP-1 logical error rate heatmaps at $p=3\times 10^{-3}$ vs. memory strength intervals. Circles mark intervals used for decoding: [0.24,0.66][-0.24,0.66] (gross), [0.161,0.815][-0.161,0.815] (two-gross), [0.254,0.985][-0.254,0.985] (surface). The dashed line indicates a threshold above which negative $\gamma_{j}$ are present. (Bottom) Relay-BP outperforms BP+OSD+CS-10 on the gross and two-gross codes, and performs comparably to Matching on the surface code. Top panels share their width scale but have separate logical error rate heat maps; bottom panels share a logical error rate scale.
@@ IV Decoding examples
[…]
In all Relay-BP simulations, we set the max iterations to $T_{r}=60$ for each leg, except the first, which uses $T_{0}=80$ due to slower initial convergence. The first leg uses a uniform memory strength of $\gamma=0.125$ for the gross and two-gross codes, and $\gamma=0.35$ for the surface code, based on preliminary sweep results. In all subsequent legs, each $\gamma_{j}$ is drawn independently and uniformly from a common range $\gamma_{\text{center}}-\gamma_{\text{width}}/2,\gamma_{\text{center}}+\gamma_{\text{width}}/2]$.
@@ V Memory strength selection
Relay-BP’s accuracy can be improved significantly by appropriate problem-dependent choice of memory strengths. Figure 2 shows heatmaps for each of the three decoding problems at $p=3\times 10^{-3}$, reporting logical error rates of Relay-BP-1 ($R=301$) with memory strengths from intervals with various values of $\gamma_{\text{center}}$ and $\gamma_{\text{width}}$. Each tile uses the smaller of 500,000 samples or a number achieving $\leq 10\%$ relative accuracy.
A striking feature in the heatmap for the gross and two-gross codes (and to a lesser extent for the surface code) is a performance hotspot above the diagonal, corresponding to intervals that include negative memory strengths. Intervals marked with circles in the heatmaps were obtained via a gradient-free optimization and are used for all subsequent simulations.
@@ VII Real-time decoding
[…]
@@
[…]
@@ Appendix B Decoding
Let $\mathbf{p}\in(0,1/2)^{N}$ be the probability vector where $p_{j}$ is the probability that error $j$ occurs. The discrete noise is then represented as an error vector $\mathbf{e}\in\mathbb{F}_{2}^{N}$, where we make the assumption that each bit is drawn independently according to a probability vector $\mathbf{p}$, such that:
@@ 9    $\hat{e}_{j}(t)\leftarrow\mathrm{HD}\bigl{(}M_{j}(t)\bigr{)}$;
10    if $\mathbf{H\hskip 0.95792pt\mathbf{\hat{e}}}(t)=\bm{\sigma}$ then